OpenStudy (anonymous):

If x+3 over 3=y+2 over 2, then x over 3=___?

OpenStudy (anonymous):

You already asked this question and I have answered it...

OpenStudy (anonymous):

(x + 3)/3 = (y + 2)/2 (x + 3) = 3(y+2)/2 x = (3(y + 2)/2) - 3 (x / 3) = ((y + 2)/2) - 1

OpenStudy (anonymous):

Lol.. Thank you !